The manager of a mens clothing store sold out of the entire stock of $60 jackets and $50 dollar jackets for a total of $3080. If the manager sold 22 more $50 dollar jackets than $60 dollar jackets. How many jackets of each were sold?
Answer by mananth(16949)

The manager of a mens clothing store sold out of the entire stock of $60 jackets and $50 dollar jackets for a total of $3080. If the manager sold 22 more $50 dollar jackets than $60 dollar jackets. How many jackets of each were sold?
Let the number of $60 jacket sold be x
So the number of $50 jackets sold will be x+22
60*x+(x+22)*50 = 3080
60x+50x+1100= 3080
110x= 3080-1100=1980
X= 18 ---- $60 jacket
Y=x=22 = 18+22= 40---- $50 jacket
mananth@hotmail.com